A group of students got the following scores in a test; 5, 10, 15, and 25. Consider samples of size 2 that can be selected from the population. What is the mean of the population?

1
Expert's answer
2022-05-11T15:31:28-0400

We have population values 5,9,10, 8, population size N=4.

Mean of population "(\mu)" = "\dfrac{5+10+15+20}{4}=12.5"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{4}(56.25+6.25+6.25+56.25)"


"=31.25"

"\sigma=\sqrt{\sigma^2}=\sqrt{31.25}\approx5.59017"